OpenStudy (anonymous):

Perform the indicated operation. Your final answer exactly as a fraction in lowest terms 11/13 - 3/15

OpenStudy (hoblos):

11/13 - 3/15 = 165/195 - 39/195 = 126/195 = 42/65